OpenStudy (anonymous):

1/2 / (Square root of 3)/2

OpenStudy (anonymous):

Like this? \[\Large \frac{\frac{1}{2}}{\frac{\sqrt{3}}{2}}\]

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

Okay... \[\frac{a}{b} \div \frac{c}{d} = \frac{a}{b} \times \frac{d}{c}\]

OpenStudy (anonymous):

I keep getting the answer 1/squr root 3

OpenStudy (anonymous):

and its supposed to be 3/ sqr root 3

OpenStudy (anonymous):

So you can re-write it as: \[\frac{1}{2} \div \frac{\sqrt{3}}{2}\] which equals \[\frac{1}{2} \times \frac{2}{\sqrt{3}}\]

OpenStudy (anonymous):

1/sqrt 3 would be correct, except you have to get rid of the squareroot in the denominator.

OpenStudy (anonymous):